# Enumeration of maps with tight boundaries and the Zhukovsky transformation

### v1

@@ 4.2 Maps with boundary-vertices
In view of Proposition 4.8 and of the relation (23) between the $T$’s and the $\tau$’s, the only task left is to show that $c(g,n)=\chi(\mathcal{M}_{g,n})$. But, as discussed in Remark 4.3, this is a consequence of [Norbury2010, Theorem 2] which asserts that the lattice count polynomial $N_{g,n}$ evaluates to $\chi(\mathcal{M}_{g,n})$ when all its variables are set to zero. ∎
@@ 5 Conclusion and discussion
In this paper, we have seen that the generating function $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ of maps of genus $g$ with $n$ tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$ appears within the expansion of $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$, a fundamental quantity of the theory of topological recursion. This coincidence is combinatorially explained by the trumpet decomposition. We have found recursion relations, also of combinatorial nature, expressing $T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}$ in terms of $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. Finally, we have seen that, when at least one $\ell_{i}$ is non-zero, $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ is equal to a parity-dependent quasi-polynomial in the variables $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ times a simple power of $R$.
At this stage, a tantalizing question is whether one can derive bijectively an expression for $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. So far, only the situation $(g,n)=(0,3)$ has been treated in [triskell] for the essentially bipartite case, and in Appendix A for the general case. In particular, the remarkably simple formulas (34) for $(g,n)=(0,4)$ and (107) for $(g,n)=(1,1)$, in the case of maps with faces of even degree only, are still waiting for a bijective proof. One may also attempt to find a combinatorial interpretation of (32), given that its constituents $p_{k}(\ell_{1},\ldots,\ell_{n})$ and $b_{n-2,k+1}$ have themselves interpretations in terms of tight maps [polytightmaps] and set partitions, respectively.
@@ Acknowledgements.
This work is supported by the Agence Nationale de la Recherche via the grants ANR-18-CE40-0033 “Dimers”, ANR-19-CE48-0011 “Combiné” and ANR-23-CE48-0018 “CartesEtPlus”. JB acknowledges the hospitality of the Laboratoire de Physique of ENS de Lyon during the initial stages of this work.
@@ Appendix A Bijective enumeration of tight pairs of pants
The purpose of this appendix is to establish the following:
@@
This theorem is an extension of [triskell, Theorem 1.1] to non bipartite maps. Note that, in this reference (which the reader is invited to consult for any definition not found in the current paper), we used the notation $T_{a,b,c}$ instead of the present notation $T^{(0)}_{2a,2b,2c}$. The (essentially) bipartite case is recovered upon taking $t_{1}=t_{3}=t_{5}=\cdots=0$ which, as noted above, implies that $R$ satisfies (4) and that $S$ vanishes. In that case, $T^{(0)}_{2a,2b,2c}$ is non-zero if and only if $a+b+c$ is an integer, as wanted.
@@
Figure 5: A summary representation of the bijection of [triskell] between maps with three tight boundaries and quintuples $(D_{1},D_{2},D_{3},T_{1},T_{2})$ made of three bigeodesic diangles $D_{1},D_{2},D_{3}$ and two bigeodesic triangles $T_{1},T_{2}$. The solid lines indicate the identification between attachment points for type I. For type II, the two outer identification lines have to be replaced by the dotted lines.
We then proceed by “conjugating” the bijective decomposition of [triskell] by $\phi$. Indeed, let $M$ be a planar map with three tight boundaries of lengths $2a,2b,2c$: clearly, $\phi(M)$ is a bipartite planar map with three tight boundaries. We may then apply one of two possible decompositions described in [ibid., Sections 5.6 and 5.7], namely:
@@
In both cases, we obtain a quintuple consisting of five pieces: two so-called bigeodesic triangles, and three bigeodesic diangles. The vertices of these pieces are those of $\phi(M)$, and we see that a midpoint vertex of $\phi(M)$ remains of degree $2$ in each piece to which it belongs: indeed, the decomposition is made by cutting along bigeodesics, and distinct bigeodesics may only merge at vertices of degree at least $3$, which are necessarily regular. As a consequence of this remark, each piece admits a (unique) preimage by $\phi$, and so our decomposition amounts to cutting $M$ itself into five pieces, denoted $(D_{1},D_{2},D_{3},T_{1},T_{2})$ on Figure 5.
@@
Figure 6: Schematic representation of a generalized bigeodesic diangle (left) and triangle (right). The tips $c_{1},c_{2},c_{3}$ are regular corners, incident to vertices (shown as solid disks). The attachment points $c_{12},...$ are either regular corners or midpoint corners, incident to midpoint vertices of the subdivided map (shown as small squares). We use the same red/blue coloring conventions as [triskell, Figures 2.2 and 2.5].
@@
With these new definitions at hand, by applying [ibid., Theorem 3.1] to $\phi(M)$, we arrive at the following:
@@
If $a\leq b+c$, $b\leq c+a$, and $c\leq a+b$, then by the type I decomposition of [triskell, Section 5.6], $M$ is in one-to-one correspondence with a quintuple $(D_{1},D_{2},D_{3},T_{1},T_{2})$ where $D_{1},D_{2},D_{3}$ are generalized bigeodesic triangles of respective exceedances $e(D_{1})=a+c-b$, $e(D_{2})=a+b-c$, $e(D_{3})=b+c-a$, and $T_{1},T_{2}$ are generalized bigeodesic triangles, with the constraints that the attachment points which are identified together on Figure 5 must be of the same (regular or midpoint) kind, and that not all elements of this quintuple are reduced to the vertex-map.
@@
Let us now discuss the enumerative consequences of Proposition A.2. To this end, it is necessary to introduce the generating functions of the various objects that may arise in the situations described above. Generally speaking, we assign a weight $t_{k}$ per inner face of degree $k$ (before subdivision), and a weight $t$ per vertex (of the map, i.e. per regular vertex of the subdivided map). Following the conventions of [ibid.], vertices that lie on strictly geodesic boundaries (displayed in red on the figures), and that are not incident to attachment points, do not receive the weight $t$. We denote by $Y$ (resp. $\widetilde{Y}$) the generating function of generalized bigeodesic triangles that are even (resp. odd). We then treat the case of diangles in the following proposition.
@@
Let $X$ denote the generating function of generalized bigeodesic diangles with exceedance 00, whose attachment points are both regular corners. Then the following holds.
@@
since $R$ (resp. $S$) counts planar maps with two distinguished boundaries both of length $1$ (resp. one of length $1$ and the other of length 00): differentiating with respect to $t$ has the effect of adding a third boundary of length 00, and boundaries of lenghts 00 or $1$ are tautologically tight. ∎
@@
Let $a$, $b$ and $c$ be nonnegative integers or half-integers, and let $T^{(0)}_{2a,2b|2c}$ (resp. $T^{(0)}_{2a|2b,2c}$) denote the generating function of planar maps with three labeled distinct tight boundaries of lengths $2a,2b,2c$, the third (resp. the second and the third) being strictly tight, where we attach a weight $t$ per vertex different from a boundary-vertex and not incident to the third boundary (resp. to the second nor to the third boundary) and, for all $k\geq 1$, a weight $t_{k}$ per inner face of degree $k$.
@@
$T^{(0)}_{2a,2b,2c}=\begin{cases}R^{2c}T^{(0)}_{2a,2b|2c}&\text{if $a+b>c$,}\\ R^{2b+2c}T^{(0)}_{2a|2b,2c}&\text{if $a>b+c$.}\\ \end{cases}$ (119)
Let us first assume $a+b>c$ and establish the formula for $T^{(0)}_{2a,2b|2c}$. If $c=0$ then there is nothing to prove, as a boundary-vertex is always considered strictly tight. Assuming $c>0$, let $M$ be a planar map with three labeled distinct tight boundaries of lengths $2a,2b,2c$. The third boundary is a face, which we denote by $F_{3}$. We then cut $M$ along the innermost minimal closed path $C$ homotopic to $F_{3}$. This path can be precisely defined as the maximal element of the lattice $\mathcal{C}_{\mathrm{min}}^{(3)}$, discussed in [triskell, Section 6.1]. We will justify in Lemma A.5 below that $C$ is simple, hence by cutting we obtain two pieces: a planar map with three tight boundaries of lengths $2a,2b,2c$, the third being strictly tight, and a planar map with two tight boundaries both of length $2c$. Moreover, the boundary resulting from this cutting operation can be canonically rooted: indeed, since planar maps with three boundaries have a trivial automorphism group, we may canonically select a distinguished corner incident to $C$. This decomposition is clearly bijective. Turning to generating functions, we decide to assign the weight $t$ for vertices on $C$ to the second map. It is known [irredmaps, Section 9.3] that planar maps with two tight boundaries both of length $2c$, one of which is rooted, have generating function $R^{2c}$. We conclude that $T^{(0)}_{2a,2b,2c}=R^{2c}T^{(0)}_{2a,2b|2c}$ as wanted.
@@
At the last step, the path $\gamma$ is reduced to $v_{0}=v_{\ell}$, and $\hat{C}$ is the sequence we are looking for. There is an interesting combinatorial structure of heap underlying this decomposition, see e.g. [Viennot1986, Proposition 6.3], but for our purposes we need only remark that the sum of the lengths of the $C_{i}$’s is $\ell$.
@@
    there exists $i,j$ such that $C_{i}$ and $C_{j}$ are homotopic to the first and second boundary, respectively.
@@
In the case where $\ell_{1}$, say, vanishes, Lemma A.5 and the above proof still hold, upon using the convention discussed in Section 1.2 and [triskell, Figure 2] of replacing the boundary-vertex by a small circle made of edges of length zero. Observe that the boundary-vertex cannot be incident to the boundary-face of length $\ell_{3}$, as otherwise following the contour of that face but circumventing the boundary-vertex in the other direction would yield a path of length $\ell_{3}$ that is homotopic to the boundary-face of length $\ell_{2}$, contradicting the tightness assumption.
The following fact, used in the proof of Lemma A.5, seems well-known in some communities, but we rederive it for convenience.
@@ B.1 A discrete integration lemma for polynomials
We first record an elementary discrete integration lemma. It is certainly well-known, but the derivation is a nice variation on the Bernoulli-Faulhaber formulas so we give a full proof.
@@
First suppose that $\ell=2n$ is even. By the Bernoulli-Faulhaber formula
@@
since the last term is the average of the two first equal terms. By a final use of the Bernoulli-Faulhaber formula, this quantity if of the form $(P(n)+P(n+1))$ where $P$ is a polynomial and $22n-1=\ell-$ 2, so it can be put in the form $P((\ell-1)/2)+P((\ell+1)/2)$, which is an even polynomial in $\ell$ by Newton’s formula. ∎
@@ B.2 Quasi-polynomiality for tight boundary insertions
where we use the simplifying notation $00_{i}=0,\ldots,$ 0 for a string of $i$ consecutive zeros. We initialize the induction at $i=0$ by noting that, by our formula (109) for tight pairs of pants,
$D_{\ell}R=R^{\frac{\ell}{2}+1}\times\begin{cases}\displaystyle{\frac{R^{\prime}}{R}}&\text{if $\ell$ is even}\\ \displaystyle{\frac{S^{\prime}}{\sqrt{R}}}&\text{ if $\ell$ is odd,}\end{cases}\qquad D_{\ell}S=R^{\frac{\ell+1}{2}}\times\begin{cases}\displaystyle{\frac{S^{\prime}}{\sqrt{R}}}&\text{if $\ell$ is even}\\ \displaystyle{\frac{R^{\prime}}{R}}&\text{ if $\ell$ is odd.}\end{cases}$ (131)
Suppose that the statement holds up to some given value of $i$: namely, we may find polynomials $P_{j,2},P_{j,3},Q_{j,1},Q_{j,2}$ in $\mathbb{Q}[x_{1},\ldots,x_{j+1},y_{1},\ldots,y_{j+1},u]$ (the second index will refer to the number of boundary-faces of odd degree in the expressions below) such that, for $0\leq j\leq i$, and letting $\epsilon$ be 00 if $\ell$ is even and $1$ if $\ell$ is odd, we have
@@
We now analyse of the right-hand side of (134) in a similar way. Its first term is equal to

### v2

@@ 4.2 Maps with boundary-vertices
In view of Proposition 4.8 and of the relation (23) between the $T$’s and the $\tau$’s, the only tasks left to show are that $c(g,n)=\chi(\mathcal{M}_{g,n})$ and that $\mathfrak{T}^{(g)}_{n}$ has total degree 3g3+n3g-3+n for every $n\geq 1$. As discussed in Remark 4.3, the first assertion is a consequence of [Nor10, Theorem 2], which states that the lattice count polynomial $N_{g,n}$ evaluates to $\chi(\mathcal{M}_{g,n})$ when all its variables are set to zero. The second assertion follows from Proposition 4.4. We believe that it should also be possible to keep track of the degree of $\mathfrak{T}^{(g)}_{n}$ in our inductive proof, at the price of a refinement of Proposition 4.8, by looking at the joint homogeneity properties of $\mathfrak{T}^{(g)}_{n}$ in the variables $\ell_{i}^{2}$ as well as the derivatives of $R$ and $S$, similarly to what is stated in Theorem C.2 for $n=0$. ∎
@@ 5 Conclusion and discussion
In this paper, we have seen that, within the theory of topological recursion applied to the enumeration of maps, the expansion of the fundamental quantity $\omega^{(g)}_{n}(z_{1},\ldots,z_{n})$ has a simple relation with the generating functions $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ of maps of genus $g$ with $n$ tight boundaries of lengths $\ell_{1},\ldots,\ell_{n}$. This coincidence is combinatorially explained by the trumpet decomposition. We have found recursion relations, also of combinatorial nature, expressing $T^{(g)}_{\ell_{1},\ldots,\ell_{n+1}}$ in terms of $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. Finally, we have seen that, when at least one $\ell_{i}$ is non-zero, $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$ is equal to a parity-dependent quasi-polynomial in the variables $\ell_{1}^{2},\ldots,\ell_{n}^{2}$ times a simple power of $R$.
At this stage, a tantalizing question is whether one can derive bijectively an expression for $T^{(g)}_{\ell_{1},\ldots,\ell_{n}}$. So far, only the situation $(g,n)=(0,3)$ has been treated in [BGM22] for the essentially bipartite case, and in Appendix A for the general case. In particular, the remarkably simple formulas (35) for $(g,n)=(0,4)$ and (110) for $(g,n)=(1,1)$, in the case of maps with faces of even degree only, are still waiting for a bijective proof. One may also attempt to find a combinatorial interpretation of (33), given that its constituents $p_{k}(\ell_{1},\ldots,\ell_{n})$ and $b_{n-2,k+1}$ have themselves interpretations in terms of tight maps [BGM24] and set partitions, respectively.
@@ Acknowledgements.
Many thanks are due to the anonymous referee for suggesting valuable improvements to this paper. This work is supported by the Agence Nationale de la Recherche via the grants ANR-18-CE40-0033 “Dimers”, ANR-19-CE48-0011 “Combiné” and ANR-23-CE48-0018 “CartesEtPlus”. JB acknowledges the hospitality of the Laboratoire de Physique of ENS de Lyon during the initial stages of this work. This work was completed while GM was holding a visiting professor position at the Research Institute for Mathematical Sciences, an International Joint Usage/Research Center located in Kyoto University.
@@ Appendix A Bijective enumeration of tight pairs of pants
The purpose of this appendix is to establish the following theorem.
@@
This theorem is an extension of [BGM22, Theorem 1.1] to non bipartite maps. Note that, in this reference (which the reader is invited to consult for any definition not found in the current paper), we used the notation $T_{a,b,c}$ instead of the present notation $T^{(0)}_{2a,2b,2c}$. The (essentially) bipartite case is recovered upon taking $t_{1}=t_{3}=t_{5}=\cdots=0$ which, as noted above, implies that $R$ satisfies (4) and that $S$ vanishes. In that case, $T^{(0)}_{2a,2b,2c}$ is non-zero if and only if $a+b+c$ is an integer, as wanted.
@@
Figure 5: A summary representation of the bijection of [BGM22] between maps with three tight boundaries and quintuples $(D_{1},D_{2},D_{3},T_{1},T_{2})$ made of three bigeodesic diangles $D_{1},D_{2},D_{3}$ and two bigeodesic triangles $T_{1},T_{2}$. The solid lines indicate the identification between attachment points for type I. For type II, the two outer identification lines have to be replaced by the dotted lines.
We then proceed by “conjugating” the bijective decomposition of [BGM22] by $\phi$. Indeed, let $M$ be a planar map with three tight boundaries of lengths $2a,2b,2c$: clearly, $\phi(M)$ is a bipartite planar map with three tight boundaries. We may then apply one of two possible decompositions described in [ibid., Sections 5.6 and 5.7], namely:
@@
In both cases, we obtain a quintuple consisting of five pieces: two so-called bigeodesic triangles, and three bigeodesic diangles. The vertices of these pieces are those of $\phi(M)$, and we see that a midpoint vertex of $\phi(M)$ remains of degree $2$ in each piece to which it belongs: indeed, the decomposition is made by cutting along bigeodesics, and distinct bigeodesics may only merge at vertices of degree at least $3$, which are necessarily regular. As a consequence of this remark, each piece admits a (unique) preimage by $\phi$, and so our decomposition amounts to cutting $M$ itself into five pieces, denoted by $(D_{1},D_{2},D_{3},T_{1},T_{2})$ in Figure 5.
@@
Figure 6: Schematic representation of a generalized bigeodesic diangle (left) and triangle (right). The tips $c_{1},c_{2},c_{3}$ are regular corners, incident to vertices (shown as solid disks). The attachment points $c_{12},$... are either regular corners or midpoint corners, incident to midpoint vertices of the subdivided map (shown as small squares). We use the same red/blue coloring convention as [BGM22, Figures 2.2 and 2.5].
@@
With these new definitions at hand, by applying [ibid., Theorem 3.1] to $\phi(M)$, we arrive at the following proposition.
@@
If $a\leq b+c$, $b\leq c+a$, and $c\leq a+b$, then by the type I decomposition of [BGM22, Section 5.6], $M$ is in one-to-one correspondence with a quintuple $(D_{1},D_{2},D_{3},T_{1},T_{2})$ where $D_{1},D_{2},D_{3}$ are generalized bigeodesic triangles of respective exceedances $e(D_{1})=a+c-b$, $e(D_{2})=a+b-c$, $e(D_{3})=b+c-a$, and $T_{1},T_{2}$ are generalized bigeodesic triangles, with the constraints that the attachment points which are identified together on Figure 5 must be of the same (regular or midpoint) kind, and that not all elements of this quintuple are reduced to the vertex-map.
@@
Let us now discuss the enumerative consequences of Proposition A.2. To this end, it is necessary to introduce the generating functions of the various objects that may arise in the situations described above. Generally speaking, we assign a weight $t_{k}$ per inner face of degree $k$ (before subdivision), and a weight $t$ per vertex (of the map, i.e. per regular vertex of the subdivided map). Following the conventions of [ibid.], vertices that lie on strictly geodesic boundaries (displayed in red in the figures), and that are not incident to attachment points, do not receive the weight $t$. We denote by $Y$ (resp. $\widetilde{Y}$) the generating function of generalized bigeodesic triangles that are even (resp. odd). We then treat the case of diangles in the following proposition.
@@
Let $X$ denote the generating function of generalized bigeodesic diangles with exceedance 0, whose attachment points are both regular corners. Then the following holds.
@@
since $R$ (resp. $S$) counts planar maps with two distinguished boundaries both of length $1$ (resp. one of length $1$ and the other of length 0): differentiating with respect to $t$ has the effect of adding a third boundary of length 0, and boundaries of lengths 0 or $1$ are tautologically tight. ∎
@@
[…]
@@
[…]
Let us first assume $a+b>c$ and establish the formula for $T^{(0)}_{2a,2b|2c}$. If $c=0$ then there is nothing to prove, as a boundary-vertex is always considered strictly tight. Assuming $c>0$, let $M$ be a planar map with three labeled distinct tight boundaries of lengths $2a,2b,2c$. The third boundary is a face, which we denote by $F_{3}$. We then cut $M$ along the innermost minimal closed path $C$ homotopic to $F_{3}$. This path can be precisely defined as the maximal element of the lattice $\mathcal{C}_{\mathrm{min}}^{(3)}$, discussed in [BGM22, Section 6.1]. We will justify in Lemma A.5 below that $C$ is simple, hence by cutting we obtain two pieces: a planar map with three tight boundaries of lengths $2a,2b,2c$, the third being strictly tight, and a planar map with two tight boundaries both of length $2c$. Moreover, the boundary resulting from this cutting operation can be canonically rooted: indeed, since planar maps with three boundaries have a trivial automorphism group, we may canonically select a distinguished corner incident to $C$. This decomposition is clearly bijective. Turning to generating functions, we decide to assign the weight $t$ for vertices on $C$ to the second map. It is known [BG14, Section 9.3] that planar maps with two tight boundaries both of length $2c$, one of which is rooted, have generating function $R^{2c}$. We conclude that $T^{(0)}_{2a,2b,2c}=R^{2c}T^{(0)}_{2a,2b|2c}$ as wanted.
@@
At the last step, the path $\gamma$ is reduced to $v_{0}=v_{\ell}$, and $\hat{C}$ is the sequence we are looking for. There is an interesting combinatorial structure of heap underlying this decomposition, see e.g. [Vie86, Proposition 6.3], but for our purposes we need only remark that the sum of the lengths of the $C_{i}$’s is $\ell$.
@@
    there exists $i,j$ such that $C_{i}$ and $C_{j}$ are homotopic to the first and second boundaries, respectively.
@@
In the case where $\ell_{1}$, say, vanishes, Lemma A.5 and the above proof still hold, upon using the convention discussed in Section 1.2 and [BGM22, Figure 2] of replacing the boundary-vertex by a small circle made of edges of length zero. Observe that the boundary-vertex cannot be incident to the boundary-face of length $\ell_{3}$, as otherwise following the contour of that face but circumventing the boundary-vertex in the other direction would yield a path of length $\ell_{3}$ that is homotopic to the boundary-face of length $\ell_{2}$, contradicting the tightness assumption.
The following fact, used in the proof of Lemma A.5, seems well known in some communities, but we rederive it for convenience.
@@ B.1 A discrete integration lemma for polynomials
We first record an elementary discrete integration lemma. It is certainly well known, but the derivation is a nice variation on the Bernoulli-Faulhaber formulas so we give a full proof.
@@
First suppose that $\ell=2n$ is even. By the Bernoulli-Faulhaber formula, we have
@@
since the last term is the average of the first two equal terms. By a final use of the Bernoulli-Faulhaber formula, this quantity if of the form $(P(n)+P(n+1))$ where $P$ is a polynomial and $2n-1=\ell-2$, so it can be put in the form $P((\ell-1)/2)+P((\ell+1)/2)$, which is an even polynomial in $\ell$ by Newton’s formula. ∎
@@ B.2 Quasi-polynomiality for tight boundary insertions
where we use the simplifying notation $0_{i}=0,\ldots,0$ for a string of $i$ consecutive zeros. We initialize the induction at $i=0$ by noting that, by our formula (112) for tight pairs of pants,
[…]
Suppose that the statement holds up to some given value of $i$: namely, we may find polynomials $P_{j,2},P_{j,3},Q_{j,1},Q_{j,2}$ in $\mathbb{Q}[x_{1},\ldots,x_{j+1},y_{1},\ldots,y_{j+1},u]$ (the second index will refer to the number of boundary-faces of odd degree in the expressions below) such that, for $0\leq j\leq i$, and letting $\epsilon$ be 0 if $\ell$ is even and $1$ if $\ell$ is odd, we have
@@
We now analyse the right-hand side of (137) in a similar way. Its first term is equal to